OpenStudy (anonymous):

***Medal AND Fan Rewarded*** How do I verify the identity ? 1 - 2cos^2x = 2sin^2x - 1

OpenStudy (nathan917):

Fraction it i would think.

OpenStudy (anonymous):

Start with the left hand side and use the fact that cos^2 = 1 - sin^2

OpenStudy (nathan917):

cos ² x - sin ² x cos ² x - (1 - cos ² x) cos ² x - 1 + cos ² x 2 cos ² x - 1 kinda like this?

OpenStudy (jack1):

sin x = cos (x-(pi/2))

OpenStudy (anonymous):

I am not getting this.....

OpenStudy (nathan917):

Pythagorean Identity and substitute it into the left hand side would help you solve this problem

OpenStudy (anonymous):

oh okay... thanks :)

OpenStudy (nathan917):

Np

OpenStudy (nathan917):

Also try this http://math.tutorvista.com/trigonometry/trigonometric-identities.html

OpenStudy (nathan917):

It should help you solve it.

OpenStudy (anonymous):

how about this, this is an identity you should know \[\sin^2x+\cos^2x=1\] multiply both sides by 2\[2\sin^2x+2\cos^2x=2\] \[2\cos^2x=2-2\sin^2x\]\[2\cos^2x-1=1-2\sin^2x\]

OpenStudy (anonymous):

I understand this now thanks and I will try out that website! Thanks for the help! @Blueskys @nathan917

OpenStudy (anonymous):

no worries

OpenStudy (nathan917):

@minnie123 Np anytime.

OpenStudy (anonymous):

New similar problem up incase your not busy and wanna take a crack at it....